OpenStudy (anonymous):

WILL GIVE MEDAL. c: The equation below shows the perimeter f(x) of several rectangular paintings having width x inches: f(x) = 40 + 2x The perimeter of each painting is less than 92 inches. Which of the following best describes the domain of f(x)? Choices: All real numbers greater than 0 and less than 26 All real numbers greater than 0 and less than 92 All integers greater than 0 and less than 26 All integers greater than 0 and less than 92

OpenStudy (mathmate):

Hint: Domain in this context means the values of x that will give 1. a positive width 2. a perimeter, f(x) less than or equal to 92. For example, if x=30, then f(x)=40+2*30=100, so 30 is not in the domain of f(x). To find the domain, you need to figure out a maximum and minimum value of x so that all values between the two extremes are valid/acceptable.

OpenStudy (anonymous):

Can you explain to me how to find the min & max?

OpenStudy (mathmate):

What is the minimum width possible?

OpenStudy (anonymous):

uhh, dunno .-. i've never done a problem like this and it's the only one on this assignment

OpenStudy (mathmate):

 How small can width be? Can it be zero? Can it be -1?

OpenStudy (anonymous):

0?

OpenStudy (mathmate):

Good, except that it cannot be exactly zero, so we say x>0.

OpenStudy (mathmate):

Now we know that the perimeter cannot exceed 92. We know the perimeter is f(x). so we look for x such that f(x)=92, that would be the maximum. Can you find that? f(x)=40+2x=92 Solve for x.

OpenStudy (anonymous):

f(x)= 40+2x=92 -40 -40 2x= 52 2x/2= 52/2 x= 26 ?

OpenStudy (mathmate):

Great! That's the maximum value x can take to make a perimeter of 92. But we were told that the perimeter must be less than 92, so what can you say about x?

OpenStudy (anonymous):

that x equals 0?

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

ugh idk I:

OpenStudy (mathmate):

We don't want f(x) to go over 92, so we limit x to be less than 26. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

So now we know that x has to be greater than zero, and less than 26. Can you put the two conditions together in a math expression?

OpenStudy (anonymous):

x<0>26?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

was i close?

OpenStudy (mathmate):

Almost! 0<x<26

OpenStudy (mathmate):

This represents the valid values of x for the given problem. Can you find the math term for this "valid values of x"?

OpenStudy (anonymous):

hmm

OpenStudy (mathmate):

How about domain!

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

x is domain and y is range XD i forgot

OpenStudy (mathmate):

Now you're ready to go and make your choice of answer! :)

OpenStudy (anonymous):

thank you :)

OpenStudy (mathmate):

*answers

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

wait is it integers or all numbers?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Good question. Since measurements of width are not limited to integers, so x can take any real number withing the extremes.